The pressure of an ideal gas is reduced by 50%, resulting in a decrease in temperature to 75% of the initial value. Calculate the ratio of the final to initial volumes of the gas.


1
Expert's answer
2021-06-21T11:38:30-0400

The combined gas law is

P1V1T1=P2V2T2\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}

Given that P2 = 0.50 P1 and T2 = 0.75 T1,

P1V1T1=0.50P1×xV10.75T1\frac{P_1V_1}{T_1}=\frac{0.50P_1\times{xV_1}}{0.75T_1}

x = 1.5, which means that V2 = 1.5 V1, and V2V1=1.5\frac{V_2}{V_1}=1.5


Answer: 1.5